# Tight contact structures on hyperbolic homology 3-spheres

### v1

@@ Abstract.
We produce a large class of hyperbolic homology $3$-spheres admitting arbitrarily many distinct tight contact structures. We also produce a sub-class admitting arbitrarily many distinct tight contact structures within the same homotopy class of oriented plane distributions. We also introduce a notion of geometric limits of contact structures compatible with geometric limits of hyperbolic manifolds. We study the behavior of the tight contact structures we construct under geometric limits.
@@
Both authors are supported in part by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500, and by an endowment of the Infosys Foundation. MM is supported in part by the Fields Institute, Toronto during the special semester on ”Randomness and Geometry”, and by a DST JC Bose Fellowship
@@ 1. Introduction
Let $M$ be an irreducible 3-manifold. Does it admit a tight contact structure? Does it admit a fillable contact structure?
The problem of existence and classification of tight contact structures on lens spaces [Hon00a] and Seifert fibered spaces [Gom98, Hon00b, LS09] is completely settled. By the work of Agol [Ago13, Theorem 9.2] and Wise [Wis21], every closed hyperbolic $3$-manifold admits a finite-sheeted cover which is fibered over the circle. Combining this with the work of Eliashberg-Thurston [ET98] on confoliations, one deduces that every closed hyperbolic $3$-manifold admits a finite-sheeted cover possessing a tight contact structure.
However, not much is known beyond this with regard to Question 1.1 for the class of hyperbolic $3$-manifolds. In this direction, there is a complete classification of tight contact structures for surgeries on the figure-eight knot, due to Conway-Min [CM20]. Moreover, there are examples of hyperbolic 3-manifolds which do not admit fillable contact structures, due to Li-Liu [LL19]. For more examples of tight contact structures on hyperbolic $3$-manifolds, see [HKM03], [AS17], [Etg12].
@@
    $M_{i,n}$ admits at least $a_{n}$ contact structures $\xi_{1},\cdots,\xi_{a_{n}}$ that are distinct up to contactomorphism,
@@
For any sequence $M_{i,n}\to M_{i,\infty}$ as in Theorem 1.2, #$\#\mathrm{Tight}(M_{n})\to\infty$ as $n\to\infty$.
@@
    In [LM97], Lisca-Matic proved an analogous result for certain Seifert-fibered spaces that are rational homology $3$-spheres, given by $1/n$-Dehn surgery on the trefoil knot.
@@
    First, we allow links $L$ in place of $K$. Such links $L$ satisfy conditions (1) and (2) above. Further, each component of $L$ satisfies condition (3).
@@
    The $1/n$-surgeries are replaced by $p/q$-surgeries (on each component of $L$). The surgery coefficients $p/q$ are required to satisfy the property that ${-q/p}$ has a terminating negative continued fraction expansion $[a_{0},\cdots,a_{n}]$ with each $a_{i}\leq-2$ (see Definition 4.1).
This allows us to construct rational homology spheres $M$ with arbitrarily many tight contact structures by rational Dehn filling on links $L$ as above (see Corollary 4.5).
@@
    $S^{3}\setminus K$ is a geometric limits of closed hyperbolic 3-manifolds equipped with contact structures,
@@
Acknowledgments: We thank Marc Kegel for comments on an earlier draft. The second author thanks Dishant Pancholi and Mike Miller Eisemeier for useful conversations.
@@ 2.1. Braid closures
denote the mapping torus of $\phi$. Fix a basepoint $z_{0}\in\partial D^{2}$. The image of $I\times\{z_{0}\}$ under the quotient map $I\times D^{2}\to T(\phi)$ gives an embedded curve $c\subset\partial T(\phi)$. Choose a genus $1$ Heegaard decomposition $S^{3}=(S^{1}\times D^{2})\cup(D^{2}\times S^{1})$, where we suppose coordinates on either torus are chosen in a way that $S^{1}\times\{z_{0}\}$ in the first solid torus bounds a meridian disk $D^{2}\times\{z_{0}\}$ in the second solid torus. Let us also choose an identification $T(\phi)\cong S^{1}\times D^{2}$ given by sending $c$ to $S^{1}\times\{z_{0}\}$. Consider the composition,
@@ 2.4. Increasing genus
    • the horizontal arcs $[p_{1},p_{2}]\times\{0\},[p_{3},p_{4}]\times\{0\}$ on $\{0\}\times D^{2}$,
    • the horizontal arcs $[p_{2},p_{3}]\times\{1\},[p_{4},p_{1}]\times\{1\}$ on $\{1\}\times D^{2}$.
@@
We summarize some conditions to ensure that a braid has hyperbolic braid closure. From Corollary 2.6 and Theorem 2.9 we have the following. Let $\beta$ be a pseudo-Anosov braid such that its closure is not one of the following:
@@ 3.1. Construction of the manifolds
$\Delta:=(\sigma_{1}\sigma_{2}\cdots\sigma_{n-1})(\sigma_{1}\sigma_{2}\cdots\sigma_{n-2})\cdots(\sigma_{1}\sigma_{2})(\sigma_{1}).$
@@
Let $M_{n}:=S_{1/n}(\widehat{\beta})$ denote the $1/n$-Dehn filling of $S^{3}\setminus\widehat{\beta}$. Here, we use the convention that $1/n$-Dehn filling attaches a disk to $\mu+n\lambda$, where $\mu$ is the meridian of the knot $\widehat{\beta}$, and $\lambda$ is a longitude given by pushing the knot off along a Seifert surface. This is an integer homology $3$-sphere, and by Thurston’s hyperbolic Dehn surgery theorem [Thu22, Chapter 4], $M_{n}$ is a closed hyperbolic $3$-manifold for $n>0$ sufficiently large.
@@ 3.2. Construction of the contact structures
Figure 2. Two Legendrian realizations of a left-handed crossing by introduction of two cusps of the same orientation.
@@ 3.3. Distinguishing contact structures
The contact structures $\xi_{n,k}$, $11\leq k\leq n-$ 1, on $M_{n}$ constructed in Section 3.2 are distinct up to isotopy.
@@
There exists a natural number $c>0$ depending only on the braid $\beta$ chosen in Section 3.1, such that for all $n>0$ sufficiently large, $\{\xi_{n,k}:1\leq k\leq n-1\}$ contains a subset of size at least $n/c$ such that all elements of the subset are pairwise distinct up to contactomorphism.
@@
By the hypothesis that $n>0$ is sufficiently large, $M_{n}$ is hyperbolic by Thurston’s Dehn surgery theorem. Suppose $\xi_{n,k}$ and $\xi_{n,l}$ are contactomorphic. As they are not isotopic by Proposition 3.6, any contactomorphism between them must be a nontrivial smooth mapping class of $M_{n}$, i.e., it must not be diffeotopic to identity. By Mostow rigidity and [GMT03], the smooth mapping class group of $M_{n}$ is isomorphic to the isometry group of $M_{n}$. If $n>0$ is chosen to be sufficiently large, then $M_{n}$ admits a unique shortest geodesic given by the core of the solid torus used in Dehn filling [Thu22]. Any isometry of $M_{n}$ must preserve this curve. Deleting this curve results in an isometry of the manifold $S^{3}\setminus K$. Therefore, $\mathrm{Isom}(M_{n})\cong\mathrm{Isom}(S^{3}\setminus K)$. Let $c:=|\mathrm{Isom}(S^{3}\setminus K)|$ denote the order of this group. This is a finite number by Mostow rigidity, and only depends on the braid $\beta$ whose closure gives the knot $K$. Thus, the action of $\mathrm{MCG}(M_{n})$ on the set $\{\xi_{n,1},\cdots,\xi_{n,n-1}\}$ has all orbits of size at most $c$. Therefore, there exist at least $n/c$ contact structures, one from each orbit, that are pairwise distinct up to contactomorphism.∎
@@ 4. Further examples: rational surgery on links
First, we set up some notation and preliminary definitions.
@@
such that $a_{i}\leq-2$. Every real number $r<-1$ admits a unique negative continued fraction expansion. For a terminating continued fraction expansion $r=[a_{0},a_{1},\cdots,a_{k}]$, we denote
$\Phi(r)=|a_{0}+1||a_{1}+1|\cdots|a_{k}+1|$
@@
Any real number with a terminating negative continued fraction expansion is rational, but the converse is not true. For instance, $-3=[-2,-2,-2,\cdots]$.
@@
where $\beta_{i}\in B_{m_{i}}$ is a braid on $m_{i}$ strands, $1\leq i\leq\ell$. We denote $c_{i,+},c_{i,-}$ to be the number of positive and negative crossings of $\beta_{i}$, respectively.
@@
Let $\beta\in B_{m}$ be a braid, with its closure $\widehat{\beta}=\widehat{\beta}_{1}\cup\cdots\cup\widehat{\beta}_{\ell}$ as above. Define a slope vector
$\mathbf{v}=(p_{1}/q_{1},\cdots,p_{\ell}/q_{\ell}),$
to be an ordered tuple of rational numbers. Let $M(\beta,\mathbf{v})$ be the $3$-manifold obtained by Dehn filling $S^{3}\setminus\widehat{\beta}$ with surgery coefficient $p_{i}/q_{i}$ on the component $\widehat{\beta}_{i}$, for all $1\leq i\leq\ell$.
Let $\mathbf{v}=(p_{1}/q_{1},\cdots,p_{\ell}/q_{\ell})$ be a slope vector such that $-q_{i}/p_{i}$ ($1\leq i\leq\ell)$ admits a terminating negative continued fraction expansion. Define
$\Phi(\mathbf{v})=\Phi(-q_{1}/p_{1})\Phi(-q_{2}/p_{2})\cdots\Phi(-q_{\ell}/p_{\ell}),$
where $\Phi(r)$ is defined in 4.1.
@@
Let $\beta\in B_{m}$ be a braid with closure $\widehat{\beta}=\widehat{\beta}_{1}\cup\cdots\cup\widehat{\beta}_{\ell}$ such that $\widehat{\beta}_{i}$ satisfies the condition $c_{i,+}-c_{i,-}-2m_{i}\geq 1$ (cf. Section 3.1) for all $1\leq i\leq\ell$. Suppose that the slope vector $\mathbf{v}=(p_{1}/q_{1},\cdots,p_{\ell}/q_{\ell})$ has the property that $-q_{i}/p_{i}$ admits a terminating negative continued fraction expansion for all $1\leq i\leq\ell$. Then $M(\beta,\mathbf{v})$ admits at least $\Phi(\mathbf{v})$ distinct tight contact structures up to isotopy.
@@
    As in Section 3.1, we construct a $4$-manifold-with-boundary $X$ such that $\partial X=M(\beta,\mathbf{v})$. This is accomplished by first writing the surgery diagram for $M(\beta,\mathbf{v})$ given by:
@@
    with surgery coefficients as the slope vector $\mathbf{v}$. We write $p_{i}/q_{i}=0-q_{i}/p_{i}$ and apply the slam-dunk operation to change surgery coefficients of $\widehat{\beta}_{i}$ to 00, at the cost of introducing a meridian unknot for $\widehat{\beta}_{i}$ with surgery coefficient $-q_{i}/p_{i}$, for all $1\leq i\leq\ell$. Then we use the terminating negative continued expansion of $-q_{i}/p_{i}$ to further modify the surgery diagram so that all the coefficients are integral, by iterated applications of slam-dunk. This replaces the meridian of each component by a chain of Hopf links, with surgery coefficients given by the continued fraction expansion. The final surgery diagram defines a Kirby diagram for the desired $4$-manifold $X$.
@@
    For each $1\leq i\leq\ell$, the conditions $c_{+,i}-c_{-,i}-2m_{i}\geq 1$ together with the conditions $a_{i}\leq-2$ (cf. Definition 4.1) for the negative continued fraction expansion for $-q_{i}/p_{i}$ allow us to modify the Kirby diagram obtained in Step $(1)$ into a Weinstein-Kirby diagram. This is accomplished by choosing a Legendrian realization of all the components (stabilizing them if necessary) and applying Theorem 3.4. Thus, we equip $X$ with the structure of a Weinstein manifold, which in turn induces a tight contact structure on $\partial X=M(\beta,\mathbf{v})$ just as in Section 3.2.
@@
    While choosing a Legendrian realization in Step $(2)$, there may be ambiguities as to the choice of the Legendrian realization of the unknots occurring in the chain of Hopf links described in Step $(1)$. For each unknot with framing $f$, the Thurston-Bennequin number is $f+1$. The possible rotation numbers of a Legendrian realization of the unknot with Thurston-Bennequin number $f+1$ are $2k-f$, where $11\leq k\leq-f-$ 1. Thus, there are $|f+1|$ such choices. As these choices can be made independently on each component of the chain of Hopf links appearing in the surgery diagram described in Step $(1)$, there are $\Phi(\mathbf{v})$ many such choices in total. The proof of Proposition 3.6 now shows that these $\Phi(\mathbf{v})$ many contact structures are all distinct up to isotopy, since the tuple of rotation numbers of all components disagree for any two such contact structures. ∎
@@
be the $n$-th convergent. Then, we observe $|p_{n}|+|q_{n}|\to\infty$, and for infinitely many $i\geq 0$, $a_{i}\leq-3$. The last assertion follows using Remark 4.2, as otherwise $r$ would be rational. Therefore, $\Phi(-q_{n}/p_{n})\to\infty$ as $n\to\infty$. Using this, we deduce the following corollary of Theorem 4.4:

### v2

@@ Abstract.
We produce a large class of hyperbolic homology $3$-spheres admitting arbitrarily many distinct tight contact structures. We also produce a sub-class admitting arbitrarily many distinct tight contact structures within the same homotopy class of oriented plane distributions. As a corollary, we give a recipe to construct hyperbolic $L$-spaces admitting arbitrarily many distinct tight contact structures. We also introduce a notion of geometric limits of contact structures compatible with geometric limits of hyperbolic manifolds and study the behavior of the tight contact structures we construct under geometric limits.
@@
Both authors are supported in part by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500, and by an endowment of the Infosys Foundation. MM was supported in part by the Fields Institute, Toronto during the special semester on ”Randomness and Geometry”, and by a DST JC Bose Fellowship
@@ 1. Introduction
Let $M$ be an irreducible 3-manifold. Does it admit a tight contact structure?
The problem of existence of tight contact structures on lens spaces [Hon00a] and Seifert fibered spaces [Gom98, Hon00b, LS09] is completely settled. By the work of Agol [Ago13, Theorem 9.2] and Wise [Wis21], every closed hyperbolic $3$-manifold admits a finite-sheeted cover which is fibered over the circle. Combining this with the work of Eliashberg-Thurston [ET98] on confoliations, one deduces that every closed hyperbolic $3$-manifold admits a finite-sheeted cover possessing a tight contact structure.
However, not much is known beyond this with regard to Question 1.1 for the class of hyperbolic $3$-manifolds. In this direction, there is a classification of tight contact structures for many surgeries on the figure-eight knot, due to Conway-Min [CM20]. Moreover, there are examples of hyperbolic 3-manifolds which do not admit fillable contact structures, independently by Kaloti-Tosun [KT17] and Li-Liu [LL19]. Also, Mark-Tosun [MT18, Theorem 1.3] produce many examples of hyperbolic manifolds with tight contact structures, by verifying that the Heegaard-Floer contact invariant does not vanish for all positive rational surgeries on certain knots, generalizing earlier works of Lisca-Stipsicz [LS04, LS11]. For more examples of tight contact structures on hyperbolic $3$-manifolds, see [HKM03], [AS17], [Etg12].
@@
    $M_{i,n}$ admits at least $a_{n}$ contact structures $\xi_{1},\cdots,\xi_{a_{n}}$ that are pairwise non-contactomorphic,
@@
For any sequence $M_{i,n}\to M_{i,\infty}$ as in Theorem 1.2, #$\#\mathrm{Tight}({M_{i,n}})\to\infty$ as $n\to\infty$.
@@
    In [LM97], Lisca-Matic proved an analogous result for certain Seifert-fibered rational homology $3$-spheres, given by $1/n$-Dehn surgery on the trefoil knot.
  2. (2)
    In [LS04], Lisca-Stipsicz proved the existence of Seifert-fibered rational homology spheres $M_{n}$ carrying at least $n$ non-contactomorphic tight, nonfillable contact structures.
@@
We also remark that Conclusion $(3)$ of Theorem 1.2 indicates that the $h$-principle, which always holds for overtwisted contact structures [Eli89], fails arbitrarily badly for tight contact structures on the manifolds that we construct.
@@
    First, we allow links $L$ in place of $K$. Such links $L$ are required to be closures of braids $\beta$ satisfying Conditions (1) and (2) above. If $\beta_{i}$, $1\leq i\leq\ell$ denotes the braids corresponding to the $i$-th cycle in the cycle decomposition of the permutation associated to $\beta$, then each $\widehat{\beta}_{i}$ is a knot and $L=\widehat{\beta}=\widehat{\beta}_{1}\cup\widehat{\beta}_{2}\cup\cdots\cup\widehat{\beta}_{\ell}$. We further require that each $\beta_{i}$ satisfy a strong form of Condition $(3)$ above. Namely, let $c_{i,+},c_{i,-}$ denote the number of positive and negative crossings of $\beta_{i}$, let $m_{i}$ denote the number of strands of $\beta_{i}$ and let $d_{i,-}$ denote the number of negative crossings between strands of $\beta_{i}$ and strands of $\beta_{j}$ for all $j\neq i$, $1\leq j\leq\ell$. Then we demand,
    $c_{i,+}-2c_{i,-}-d_{i,-}-m_{i}\geq 1.$
@@
    The $1/n$-surgeries are replaced by positive $p/q$-surgeries (on each component of $L$). This is done in two stages, with $p/q\in(0,1)$ in Section 4.1 and then with arbitrary $p/q>0$ in Section 4.2. The key observation in the $p/q\in(0,1)$ case is that ${-q/p}$ has a terminating negative continued fraction expansion $[a_{0},\cdots,a_{n}]$ with each $a_{i}\leq-2$ (see Definition 4.1). For the general $p/q>0$ case, we effectively reduce to the $p/q\in(0,1)$ case by iterated applications of the slam-dunk operation on multiple disjoint copies of the unknotted meridian, see Proposition 4.9, and Corollaries 4.11 and 4.13.
This allows us to construct hyperbolic rational homology spheres $M$ with arbitrarily many tight contact structures by rational Dehn filling on links $L$ as above (see Corollary 4.8). Moreover, we also obtain a large class of hyperbolic knots and links, such that all positive rational surgeries on them admit tight (in fact, Weinstein-fillable) contact structures (see Corollary 4.11 and Corollary 4.13). We use this result in Section 4.3 to give a recipe for producing hyperbolic $L$-spaces $M_{n}$ with #$\#\mathrm{Tight}(M_{n})\to\infty$ (see Proposition 4.14 and Proposition 4.15, as well as Example 4.16).
@@
    $S^{3}\setminus K$ is a geometric limit of closed hyperbolic 3-manifolds equipped with contact structures,
@@
Acknowledgments: We thank Marc Kegel for comments on an earlier draft, and Bülent Tosun for pointing us to the references [MT18] and [KT17]. The second author thanks Dishant Pancholi and Mike Miller Eisemeier for useful conversations.
@@ 2.1. Braid closures
denote the mapping torus of $\phi$. Fix a base-point $z_{0}\in\partial D^{2}$. The image of $I\times\{z_{0}\}$ under the quotient map $I\times D^{2}\to T(\phi)$ gives an embedded curve $c\subset\partial T(\phi)$. Choose a genus $1$ Heegaard decomposition $S^{3}=(S^{1}\times D^{2})\cup(D^{2}\times S^{1})$, where we suppose coordinates on either torus are chosen in a way that $S^{1}\times\{z_{0}\}$ in the first solid torus bounds a meridian disk $D^{2}\times\{z_{0}\}$ in the second solid torus. Let us also choose an identification $T(\phi)\cong S^{1}\times D^{2}$ given by sending $c$ to $S^{1}\times\{z_{0}\}$. Consider the composition,
@@ 2.4. Increasing genus
    • the horizontal arcs $\{0\}\times[p_{1},p_{2}],\{0\}\times[p_{3},p_{4}]$ on $\{0\}\times D^{2}$,
    • the horizontal arcs $\{1\}\times[p_{2},p_{3}]$, $\{1\}\times[p_{4},p_{1}]$ on $\{1\}\times D^{2}$.
@@
We summarize some conditions to ensure that a braid has hyperbolic braid closure. From Corollary 2.6 and Theorem 2.9 we have the following. Let $\beta$ be a pseudo-Anosov braid. Then $\widehat{\beta^{k}}$ is hyperbolic for $k>4$, provided $\widehat{\beta^{k}}$ is not one of the following:
@@ 3.1. Construction of the manifolds
$\Delta:=(\sigma_{1}\sigma_{2}\cdots\sigma_{m-1})(\sigma_{1}\sigma_{2}\cdots\sigma_{m-2})\cdots(\sigma_{1}\sigma_{2})(\sigma_{1}).$
@@
Let $M_{n}=S^{3}_{1/n}(\widehat{\beta})$ denote the $1/n$-Dehn filling of $S^{3}\setminus\widehat{\beta}$. Here, we use the convention that $1/n$-Dehn filling attaches a disk to $\mu+n\lambda$, where $\mu$ is the meridian of the knot $\widehat{\beta}$, and $\lambda$ is a longitude given by pushing the knot off along a Seifert surface. This is an integer homology $3$-sphere, and by Thurston’s hyperbolic Dehn surgery theorem [Thu22, Chapter 4], $M_{n}$ is a closed hyperbolic $3$-manifold for $n>0$ sufficiently large.
@@ 3.2. Construction of the contact structures
Figure 3. Two possible Legendrian realizations of a left-handed crossing by introduction of two cusps of the same orientation.
@@ 3.3. Distinguishing contact structures
The contact structures $\xi_{n,k}$, $11\leq k\leq n-$ 1, on $M_{n}$ constructed in Section 3.2 are pairwise non-isotopic.
@@
There exists a natural number $c>0$ depending only on the braid $\beta$ chosen in Section 3.1, such that for all $n>0$ sufficiently large, $\{\xi_{n,k}:1\leq k\leq n-1\}$ contains a subset of size at least $n/c$ such that all elements of the subset are pairwise non-contactomorphic.
@@
By the hypothesis that $n>0$ is sufficiently large, $M_{n}$ is hyperbolic by Thurston’s Dehn surgery theorem. Suppose $\xi_{n,k}$ and $\xi_{n,l}$ are contactomorphic. As they are not isotopic by Proposition 3.6, any contactomorphism between them must be a nontrivial smooth mapping class of $M_{n}$, i.e., it must not be diffeotopic to identity. By Mostow rigidity and [GMT03], the smooth mapping class group of $M_{n}$ is isomorphic to the isometry group of $M_{n}$. If $n>0$ is chosen to be sufficiently large, then $M_{n}$ admits a unique shortest geodesic given by the core of the solid torus used in Dehn filling [Thu22]. Any isometry of $M_{n}$ must preserve this curve. Deleting this curve results in an isometry of the manifold $S^{3}\setminus K$. Therefore, $\mathrm{Isom}(M_{n})\cong\mathrm{Isom}(S^{3}\setminus K)$. Let $c:=|\mathrm{Isom}(S^{3}\setminus K)|$ denote the order of this group. This is a finite number by Mostow rigidity, and only depends on the braid $\beta$ whose closure gives the knot $K$. Thus, the action of $\mathrm{MCG}(M_{n})$ on the set $\{\xi_{n,1},\cdots,\xi_{n,n-1}\}$ has all orbits of size at most $c$. Therefore, there exist at least $n/c$ contact structures, one from each orbit, that are pairwise non-contactomorphic.∎
@@ 4. Further examples: rational surgery on links
We begin with some notation and preliminary definitions.
@@
such that $a_{i}\leq-2$. These are sometimes known in the literature as Hirzebruch-Jung continued fractions.
@@
We make a few remarks on the existence and termination of negative continued fraction expansions (Definition 4.1).
  1. (1)
    Every real number $r<-1$ admits a negative continued fraction expansion. We briefly describe an algorithm to produce such a thing. Set $a_{0}:=-\lceil-r\rceil$. Note $a_{0}\leq-2$. If $r$ is an integer, the algorithm terminates and $r=[a_{0}]$ is the required negative continued fraction expansion. Otherwise, observe:
    $r=a_{0}-\frac{1}{(a_{0}-r)^{-1}}$
    Now, $a_{0}-r=-\lceil-r\rceil-r\in(0,1)$, as $r$ is not an integer. Thus, $(a_{0}-r)^{-1}$ is well-defined and less than $-1$. Reassign $r$ to be $(a_{0}-r)^{-1}$ above, and proceed in the same fashion to define $a_{1},a_{2},\cdots$. Note that we terminate upon encountering an integer in the process. We also have $a_{i}\leq-2$. Then, $r=[a_{0},a_{1},a_{2},\cdots]$ is a negative continued fraction expansion.
  2. (2)
    Any real number with a terminating negative continued fraction expansion is rational, but the converse is false. For instance, $-1=[-2,-2,-2,\cdots]$.
    However, if $r<-1$ is a rational number, then the algorithm above does produce a terminating negative continued fraction for $r$. To see this, suppose $r=-m/n<-1$, where $m/n>0$ and $(m,n)=1$. The above algorithm can be described as a variant of the Euclidean algorithm. Indeed, find the unique integer $d$ such that $m=nd-k$, where $10\leq k\leq n-$ 1. Then $a_{0}=-d$. If $k=0$, the algorithm terminates. If not, we repeat the process with $(a_{0}-r)^{-1}=-n/k$. We then find the unique integer $d^{\prime}$ such that $n=kd^{\prime}-l$, where $10\leq l\leq k-$ 1. Notice we have a strict inequality $l<k$. This means eventually the algorithm must terminate. Hence, we have produced a terminating continued fraction expansion for $r$.
Definition 4.3.
Let $r<-1$ be a real number. Suppose the algorithm described in Remark 4.2 produces a terminating continued fraction expansion $r=[a_{0},a_{1},\cdots,a_{k}]$. Then, we shall denote
$\Phi(r):=|a_{0}+1||a_{1}+1|\cdots|a_{k}+1|$
4.1. Rational surgery on links: $10<p/q<$ 1
@@
where $\beta_{i}\in B_{m_{i}}$ is a braid on $m_{i}$ strands, $1\leq i\leq\ell$. We fix the following notation:
  1. (1)
    Let $c_{i,+},c_{i,-}$ be the number of positive and negative crossings of $\beta_{i}$, respectively.
  2. (2)
    Let $d_{i,j,-}$ be the number of negative crossings between strands of $\beta_{i}$ and strands of $\beta_{j}$. We also define,
    $d_{i,-}=\sum_{j\neq i}d_{i,j,-}$
Remark 4.4.
While the bookkeeping of $c_{i,-}$ plays a crucial role in ensuring an appropriate Legendrian realization of each component $\widehat{\beta}_{i}$ of $\widehat{\beta}$, an analogous role is played by the quantity $d_{i,-}$ in ensuring an appropriate Legendrian realization of the crossings between different components $\widehat{\beta}_{i}$ and $\widehat{\beta}_{j}$ of $\widehat{\beta}$. We shall see this in detail in the proof of Theorem 4.7.
@@
Let $\beta\in B_{m}$ be a braid, with its closure $\widehat{\beta}=\widehat{\beta}_{1}\cup\cdots\cup\widehat{\beta}_{\ell}$ as above. Define a slope vector to be an ordered tuple of rational numbers,
$\mathbf{v}=(p_{1}/q_{1},\cdots,p_{\ell}/q_{\ell}).$
Let $M(\beta,\mathbf{v})$ be the $3$-manifold obtained by Dehn filling $S^{3}\setminus\widehat{\beta}$ with surgery coefficient $p_{i}/q_{i}$ on the component $\widehat{\beta}_{i}$, for all $1\leq i\leq\ell$.
Definition 4.6.
Let $\mathbf{v}=(p_{1}/q_{1},\cdots,p_{\ell}/q_{\ell})$ be a slope vector such that $p_{i}/q_{i}\in(0,1)$ ($1\leq i\leq\ell)$. By Remark 4.2, $-q_{i}/p_{i}$ has a terminating negative continued fraction expansion. We define
$\Phi(\mathbf{v}):=\Phi(-q_{1}/p_{1})\Phi(-q_{2}/p_{2})\cdots\Phi(-q_{\ell}/p_{\ell}),$
where $\Phi(r)$ is defined in Definition 4.1.
@@
Let $\beta\in B_{m}$ be a braid with closure $\widehat{\beta}=\widehat{\beta}_{1}\cup\cdots\cup\widehat{\beta}_{\ell}$ such that ${\beta}_{i}$ satisfies the condition $c_{i,+}-2c_{i,-}-d_{i,-}-m_{i}\geq 1$ for all $1\leq i\leq\ell$. Let $\mathbf{v}=(p_{1}/q_{1},\cdots,p_{\ell}/q_{\ell})$ be a slope vector such that $10<p_{i}/q_{i}<$ 1 for all $1\leq i\leq\ell$. Then $M(\beta,\mathbf{v})$ admits at least $\Phi(\mathbf{v})$ distinct tight contact structures up to isotopy.
@@
    As in Section 3.1, we construct a smooth $4$-manifold-with-boundary $X$ such that $\partial X=M(\beta,\mathbf{v})$. This is accomplished by first writing the smooth surgery diagram for $M(\beta,\mathbf{v})$ given by:
@@
    with surgery coefficients as the slope vector $\mathbf{v}$. We write $p_{i}/q_{i}=0-q_{i}/p_{i}$ and apply the slam-dunk operation to change surgery coefficients of $\widehat{\beta}_{i}$ to 00, at the cost of introducing an unknotted meridian for $\widehat{\beta}_{i}$ with surgery coefficient $-q_{i}/p_{i}$, for all $1\leq i\leq\ell$. We express $-q_{i}/p_{i}$ as a terminating negative continued fraction expansion:
    $-q_{i}/p_{i}=[a_{i,0},a_{i,1},\cdots,a_{i,k_{i}}].$
    We further modify the surgery diagram so that all the coefficients are integral, by iterated applications of slam-dunk. Namely, replace the meridian of each component $\widehat{\beta_{i}}$ with surgery coefficient $-q_{i}/p_{i}$, by a chain of Hopf links with surgery coefficients $a_{i,0},a_{i,1},\cdots,a_{i,k_{i}}$ as in the continued fraction expansion. The final framed link defines a Kirby diagram for the desired smooth $4$-manifold $X$.
@@
    Next, we upgrade the smooth Kirby diagram in Step $(1)$ to a Weinstein-Kirby diagram. This entails choosing a Legendrian representative of the framed link at the end of Step $(1)$, so that the framing condition “$f=\mathrm{tb}-1$” in Theorem 3.4 holds.
    To do this, we first use the technique from the proof of Theorem 3.3 to choose a Legendrian representative for each component $\widehat{\beta}_{i}$ with Thurston-Bennequin number $c_{i,+}-2c_{i,-}-m_{i}$, for all $1\leq i\leq\ell$. Then, for any pair of negatively crossing strands between $\beta_{i}$ and $\beta_{j}$, we choose a Legendrian realization as in any one of the two pictures in Figure 3. This procedure introduces a pair of cusps on (the Legendrian realization of) either $\widehat{\beta}_{i}$ or $\widehat{\beta}_{j}$. Thus, for each $1\leq i\leq\ell$, this procedure introduces at most $2d_{i,-}$ cusps on the Legendrian realization of $\widehat{\beta}_{i}$. Therefore, at the end of the procedure, the Thurston-Bennequin number for $\widehat{\beta}_{i}$ is at least:
    $(c_{i,+}-2c_{i,-}-m_{i})-\frac{1}{2}\cdot 2d_{i,-}=c_{i,+}-2c_{i,-}-d_{i,-}-m_{i},$
    which is at least $1$ by hypothesis. We apply some number of stabilizations to ensure that the Thurston-Bennequin number is $1$, for all $1\leq i\leq\ell$.
@@
    Finally, for each of the unknots occurring in the chain of Hopf links described in the framed link in Step $(1)$, we choose a Legendrian realization. This is done so that for any of the unknotted components occurring with framing $f$, the Thurston-Bennequin number is $f+1$. This is possible, since $f$ occurs as one of the terms in the negative continued fraction expansion of $-q_{i}/p_{i}$ for some $1\leq i\leq\ell$. Therefore, by definition, $f\leq-2$ (see Definition 4.1).
    Observe that the possible rotation numbers of a Legendrian unknot with Thurston-Bennequin number $f+1$ are $2k-f$, $11\leq k\leq-f-$ 1. Thus, there are $|f+1|$ such choices. As these choices can be made independently on the rest of the components, there are at least $\Phi(\mathbf{v})$ many such choices in constructing the required Legendrian realization of the entire link.
  4. (4)
    Step $(2)$ and Step $(3)$ constructs at least $\Phi(\mathbf{v})$ distinct Legendrian realizations of the smooth framed link defining the smooth $4$-manifold $X$ in Step $(1)$. Thus, by Theorem 3.4, we obtain at least $\Phi(\mathbf{v})$ potentially distinct Weinstein structures on $X$. In turn, that gives at least $\Phi(\mathbf{v})$ potentially distinct tight contact structures on $\partial X=M(\beta,\mathbf{v})$. The proof of Proposition 3.6 now shows that these $\Phi(\mathbf{v})$ many contact structures are indeed pairwise-nonisotopic, since the tuple of rotation numbers of all the components disagree for any two such contact structures. This concludes the proof. ∎
@@
be the $n$-th convergent. Then, we observe $|p_{n}|+|q_{n}|\to\infty$, and for infinitely many $i\geq 0$, $a_{i}\leq-3$. The last assertion follows from the observation $-1=[-2,-2,-2,\cdots]$ in Remark 4.2, as otherwise $r$ would be rational. Therefore, $\Phi(-q_{n}/p_{n})\to\infty$ as $n\to\infty$. Using this, we deduce the following corollary of Theorem 4.7: